OpenStudy (anonymous):

What is the sum? 3/4 + 5/16

jimthompson5910 (jim_thompson5910):

the denominators (4 and 16) are NOT the same, so you cannot add the numerators straight away what you have to do is get the denominators both equal to some number. So multiply top and bottom of 3/4 by 4 to get 12/16 therefore 3/4 + 5/16 = 12/16 + 5/16 3/4 + 5/16 = (12+5)/16 3/4 + 5/16 = 17/16 and you can decide to leave it as an improper fraction or convert to a mixed number

OpenStudy (anonymous):

I got it thanks. 1 1/16

jimthompson5910 (jim_thompson5910):

yep, \[\Large \frac{17}{16} = 1 \frac{1}{16}\]